OpenStudy (anonymous):

Find the value of (f ◦ g)' at the given value of x. f(u)=1/(cos^4 u) -1 , u=g(x)=pi x , x=10 I got -1 as an answer but its not an answer choice

OpenStudy (anonymous):

Ok. So (f o g)' is taking derivative of function over function. \[f(u) = \frac{1}{\cos^4 u} -1 ; u = g(x) = \pi x\] \[f \prime = -4 \cos^{-5} u (-sinu) u \prime = \frac{4sinu}{\cos^5u} \pi\] Substitute u = 10pi then we obtain \[f \prime = \frac{4\pi \sin10\pi}{\cos^5 {10\pi}}\] But \[\sin (10\pi) = \sin (0 + 5.2\pi) = \sin (0) = 0\] Then the answer is 0

OpenStudy (anonymous):

thats not an answer choice. they are: -pi -10pi 10pi 4-pi

OpenStudy (anonymous):

Could you please use equation editor to type f(u) again. I am confused between \[f(u) = \frac{1}{\cos^4 u} - 1\] or \[f(u) = \frac{1}{\cos^4 u - 1}\]

OpenStudy (anonymous):

the first one you wrote down. im on mobile so its hard to do the equation writer.

OpenStudy (anonymous):

I cannot find another solution, please check your question again. Actually, the method is taking the derivative of function f over function g. Then I did it, with sin 10pi = 0, we have the answer should be 0